OpenStudy (anonymous):

given sin theta =-3/5 and csc theta =-5/3 in quadrant 3, how do I find the value of other trigonometric functions using a Pythagorean identity?

Nnesha (nnesha):

 make a triangle attt 3rd quadrant \[\rm sin = \frac{ opposite }{ hypotenuse }~~~~ \cos = \frac{ adjacent }{ hypotenuse } ~~\tan= \frac{ opposite }{ adjacent }\]

Nnesha (nnesha):

 sin is opposite over hypotenuse so apply Pythagorean theorem to find 3rd side of right triangle a^2+b^2 =c^2